A satellite orbiting the earth in a circular orbit of radius R completes 1 revolution in 3 hrs orbital radius of geostationary satellite is 42000 km the orbital radius R is a,7000 b,10500 c,14000 d,1500p

Grade:11

1 Answers

Arun
6 years ago
 
Dear Tushar
 
(T1/T2)² = (R1/R2)³
Here T1 = 3 hr, T2 = 24 hr., R2 = 42000 km
Hence
(1/64 )1/3= R1/42000
R1 = 42000/4
R1 = 10500 km
 
Regards
Arun (askIITians forum expert)